For certain group of individuals, the average cost of a trip to the super bowl was $875. The standard deviation for the population was $50. This year, 20 fans who scheduled the trip paid an average of $890. Test the claim that the average cost has gone up to attend the super bowl. Use a 0.04 level of significance?
Answer by stanbon(48568)

For certain group of individuals, the average cost of a trip to the super bowl was $875. The standard deviation for the population was $50. This year, 20 fans who scheduled the trip paid an average of $890. Test the claim that the average cost has gone up to attend the super bowl. Use a 0.04 level of significance?
-------------------------
Ho: u = 875
Ho: u > 875
--------------------------------
z(890) = (890-875)/[50/sqrt(20)] = 1.3416
p-value = P(t>1.3416 with df=19) = 0.0185
------------
Conclusion:
Since the p-value is less than 4%, reject Ho.
The test results support the claim that the average cost is up.
=============================================================
Cheers,
Stan H.